OpenStudy (anonymous):

Find exact value cos(pi/4) - sin(pi/6)

OpenStudy (austinl):

Use this!

OpenStudy (anonymous):

I got (2-\[\sqrt{2}\])/2\[\sqrt{2}\]

OpenStudy (austinl):

What?

OpenStudy (anonymous):

and its still wrong (2-Sqrt2)/(2sqrt2)

OpenStudy (anonymous):

Can it be simplified??

OpenStudy (austinl):

Using the Unit Circle we get \[cos(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\] \[sin(\frac{\pi}{6})=\frac{1}{2}\]

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{2}}{2}- \cfrac{\sqrt{3}}{2}\)

OpenStudy (jdoe0001):

ohh shoot is sine my bad

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{2}}{2}- \cfrac{1}{2}\)

OpenStudy (anonymous):

i got that far

OpenStudy (jdoe0001):

get the lcd, add the rationals, all you can simplify really

OpenStudy (austinl):

So from that we are given, \[\frac{\sqrt{2}}{2}-\frac{1}{2}\]=\[\frac{\sqrt{2}-1}{2}\]

OpenStudy (anonymous):

One word Awesome!!!